OpenStudy (anonymous):

write the equation of a line parallel to the line 2x+3y=5 and passes through the point (-2,1) A) 2x-3y=7 B) 3x-2y=7 C) 3x+2y=5 D) 2x+37=-1

OpenStudy (ranga):

Put the given equation in slope-intercept form and find the slope. Parallel lines have the same slope. Plug the slope value m in: y = mx + b Since it passes through (-2,1), put x = -2 and y = 1 and solve for b.

OpenStudy (anonymous):

I got y=2x+5, but it doesn't appear to be in the right form. Could you help me on converting it?

OpenStudy (ranga):

2x+3y=5 3y = -2x + 5 y = -2/3x + 5/3 slope = -2/3 Parallel line will have the same slope. y = -2/3x + b passes through (-2,1) 1 = -2/3(-2) + b 1 = 4/3 + b b = 1-4/3 = -1/3 y = -2/3x - 1/3 multiply by 3 3y = -2x - 1 2x + 3y = -1